You are pulling a child on a sled across a frozen pond (the ice makes it frictionless). The mass of the child-plus-sled is 20kg, and you exert a force of 40N for 3 seconds, starting from rest.

(a) What is the acceleration of the sled?
(b) How far does the sled go?
(c) How fast is it going, at the end?
(d) How much work did you do?
(e) What is the final kinetic energy of the sled-plus-child?
(f) What was your power output?
Answer #1

A) ACCELERATION = net force / mass

              = 40/20 = 2 m/s2

B) at the beginning it moves with 0 velocity

C) at the end it moves with

v = u + at

v = 0 + 2*3 = 6 m/s

D) work done = change in Kinetic energy

W = 0.5mv^2 = 0.5*20*6^2 = 360 J

e) initial KE = 0

final KE = 360 - 0 = 360 J

e) power =work done / time = 360/3= 120W